### v2

# Address-addressable read AI artificial intelligence

Abstract
The graph database (GDB) is an increasingly common storage model for data involving relationships between entries. Beyond its widespread usage in database industries, the advantages of GDBs indicate a strong potential in constructing symbolic artificial intelligences and retrieval-augmented generation (RAG), where knowledge of data inter-relationships takes a critical role in implementation. However, current GDB models are not optimised for hardware acceleration, leading to bottlenecks in storage capacity and computational efficiency. In this paper, we propose a hardware-friendly GDB model, called Views. We show its data structure and organisation tailored for efficient storage and retrieval of graph data and demonstrate its functional equivalence and storage performance advantage compared to represent traditional graph representations. We further demonstrate its symbolic processing abilities in semantic reasoning and cognitive modelling with practical examples and provide a short perspective on future developments.
Subjects – artificial intelligence, electrical engineering
Keywords – graph database, knowledge representation, near-memory computing, semantic processing
1 Introduction
GDBs are used to describe, organise and manipulate data in the form of graphs, while also using graph-based integrity checks. This endorses efficient storage/retrieval and visualisation of interrelated data, especially when the workload requires frequent “traversals” of the graph, i.e., the sequential retrieval of related (a.k.a. “connected”) data. GDBs are frequently seen in areas where data connections have the same or higher importance as the data, from industrial data analysis (consider a graph detailing the supplier/client relationships in a supply chain) to the combination of deep learning with neuromorphic research (for example so-called “RAG pipelines”).
Widespread usage of GDB is seen in knowledge representations, particularly in implementing knowledge graphs – most commonly instantiated as Resource Description Framework (RDF) triplestores or label property graphs[1]. RDF represents knowledge as “subject-predicate-object” triples with optional reification and named graphs for a clean logical footing [2, 3]. LPGs, on the other hand, attach labels and properties directly to vertices and edges for application-centric schemas and analytics. Across both families, the explicitness of knowledge graph semantics shows promise in symbolic cognitive computing by exposing symbols and relations to reasoning procedures [4, 5]. Knowledge graphs are also increasingly important as knowledge bases for large language models in RAG pipelines to improve their accuracy and reduce hallucinations [6]. These then call for more efficient graph data storage and processing methods to handle the ever-growing size and complexity of knowledge graphs.
Operationally, querying is the critical workload in GDBs. For example, the query of “find me all films which Tom Hanks has acted in” followed by “who is Tom Hanks”, as anyone might do in any film database. Multiple query languages and formal frameworks have been proposed over the years for expressive graph pattern matching and path traversal in RDF and LPGs, such as GQL, SPARQL, Cypher, etc. However, these high-level languages are built upon but not necessarily optimised for existing computing facilities [7, 8]. To address these problems, contemporary solutions such as GPU clusters can alleviate some GDB workload throughput limits, although the high cost of power and deployment complexity constrain their widespread adoption [9].
As a result, contemporary GDB models and frameworks face significant challenges. In symbol-heavy applications like knowledge graphs, where content search operations dominate, this manifests as poor cache behaviour and bandwidth pressure. Heterogeneity of data types and cross-database operability further present a bottleneck in efficient processing and scaling [10, 11], for example in the domain of biological science[12]. Memory and schema optimisations are regarded as promising ways to enhance storage efficiency and query performance and to alleviate such issues [12, 13, 14], while practical hardware implementation considerations are still overlooked.
Now we are going to talk about what to do in cases of data of unusual complexity. Historically, recursive labelling in GDB models has been proposed as a method to represent complex relationships in graphs, especially in semantic contexts. Pratt developed the initial hierarchical graph model of recursive labelling for semantics of programming languages [15], and Boley proposed directed recursive labelnode hypergraph (DRLH) as a representation language for semantic network representations, capable of being specialised into natural language or other languages such as Lisp [16]. However, these methods are not directly applicable to modern GDBs, due to the lack of support for graph operation practices and operational methods for database-level processing. They did not develop graph models further to a direct hardware implementation either.
As an overview, current GDB scalability challenges call for the development of specialised hardware accelerators that: a) can “natively” store information in graph format and b) can use in- and/or near-memory computation to perform graph-oriented operations on said data in a massively parallel fashion. In other words, we need to develop a data structure tailor-made for GDBs and an accompanying hardware that will store and operate on it. By processing graph data on such dedicated hardware, storage efficiency enhancement and direct real-time manipulations become feasible, leading to higher operation speed.
In this work, we propose Views, a graph database model tailored for use in hardware accelerators. First, we show how at its foundation it refactors descriptions of directed and labelled graphs as linked lists for better uniformity and symmetry in the data structure. Then we proceed to specify the exact data organisation structure that enables a piece of physical hardware to implement the GDB model and show both how it can represent a labelled graph and resolve semantic queries. Finally, we provide a few toy examples of how this GDB model could be used to perform simple “cognitive” tasks on a small dataset consisting of letters and strings.
This paper is organised as follows: Section 2 describes the proposed Views GDB model in terms of its data encoding and structure, its equivalence to conventional graph representations, and its instantiation on graph database hardware. Section 3 presents the hardware implementation methodology with a comparison between existing graph database implementations, outlining 2 mapping schemes and Associative Chip Architecture (ASOCA) project for hardware acceleration. Section 4 demonstrates the operation on Views through examples in semantic reasoning and a Copycat-inspired cognitive processing application [17]. Section 5 provides a high-level discussion on the relationship of Views to other representations and potential extensions.
2 The Proposed Views Model
The proposed Views GDB model has been designed to represent graphs with the following features: (1) Directedness: vertex pairs are connected by directed edges (or “arcs”), (2) Labellability: edges and vertices can both be “labelled”, i.e., we can attach further properties to them and (3) Recursive properties: properties themselves can have properties and so on ad infinitum, rendering the capability of representing complex, nested relationships. We call these directed recursively labelled graphs, a term inspired by Boley’s DRLH [16] but restricting the underlying structure to graphs rather than hypergraphs, and any mention of “graph” in the rest of this paper will refer to a DRLG unless otherwise stated.
We shall now introduce the basic data structure of the proposed GDB model and then explain its ability to support infinitely recursive labellability as well as its mapping to hardware. Finally, the building and organisation of a Views-based GDB is presented.
[…]
sourcedestinationlabel
sourcedestinationlabel
Figure 1: Forms of basic data structure: (a) A “half-$K_{2}$” graph. Rectangles represent abstract graph vertices and the arrow represents an abstract graph edge. (b) The triplet in Views GDB model. Here, the bevelled rectangles represent data that can be stored in physical memory entries as numbers or pointers.
We begin construction of the Views data structure by considering a simple, non-trivial DRLG as illustrated in Figure 1a, which consists of two vertices (denoted in non-bevelled rectangles) and a single, labelled, directed edge connecting them. We call this a “half-$K_{2}$” digraph (to distinguish it from the full $K_{2}$ digraph which includes the reciprocal edge) and treat it as the “unit of relational information” within a graph: repeated instantiation of such ternary relationships can be used to create any DRLG. In the semantic web context, this corresponds to a RDF triple, and repeated instantiation constructs an RDF graph [2, 3]. Within the form of RDF triple, ¡subject-predicate-object¿, subject corresponds to the “source” vertex, predicate to the labelled edge, and object to the “destination” vertex. Finally, note that the terms “source”, “edge” and “destination” are used in a “de re” fashion and refer to the objects in question directly: we have not yet made use of labellability so far.
Views represents the conjugate ternary relationship using a “source-centred”, numerical structure that we shall call the Views triplet, meaning that edges and destination vertices are treated as equivalent entities (i.e., we make no distinction between them), both of which relate to the source vertex in the same fashion (as indicated by the “-t’ suffix’s shape). This is illustrated in Figure 1b, where the source vertex “points to” both the edge and the destination vertex. Note that bevelled rectangles in the discourse differ from conventional vertices and they refer to physical memory entries as numbers or pointers, and that the arrows in this figure denote the concept of “owns”, i.e., the source vertex “owns” the marked destination vertex and edge. Thus, the pair of vertices and connecting edge have been mapped onto a set of 3 numbers that can be stored in a physical memory. So long as there is a way to track the association between the members of the Views triplet, we can represent a triplet in physical memory (more details in subsequent sections).
2.2 From Triplet to “Simple” Vertex-labelled GDB
It is easy to see that multiple instances of the numerical triplet structure above can collectively form a GDB. To begin building a GDB with this structure we begin by adding multiple edges to some source vertex and to do this in practice we upgrade the Views triplet from above into a simplified form of the data structure that Views relies upon (the “linknode”): the proto-linknode. This “proto-linknode” stores the unlabelled triplet information from the previous section, mapping it to a set of three identifier numbers. Internally, a proto-linknode is formatted as follows: [“source vertex”, “edge”, “destination vertex”, “next linknode”], abbreviated to: [head ID, primID1, primID2, next], where primID stands for “primary ID”, illustrated in Figure 2a. head ID corresponds to the “source vertex” whilst primID1 and primID2 correspond to the “edge” and “destination vertex” respectively. Note how the updated naming convention makes it explicitly clear that there is no distinction in principle between edge and destination vertex.
link addresshead IDprimID1primID2next
link addresshead IDNULLNULLnext
Figure 2: Views GDB model (a) proto-linknode, and (b) proto-headnode. Ellipses in this figure refer to the physical addresses corresponding to each portrayed link/headnode, but are not explicitly stored in each linknode’s allocated memory space. Bevelled rectangles are data explicitly stored in the linknode’s allocated memory. Therefore, link address diagrammatically represents the address of the current linknode and next is a physically stored pointer to the next linknode addresses while head ID in red is another physically stored pointer to the source vertex. Note that in the case of a headnode, head ID stores the same value as link address i.e. it points to itself. This also relates back to Figure 1b.
With the linknode partly defined we can now add multiple linknodes with the same head ID in order to add further properties to some source vertex; each corresponds to an outgoing edge with its label and destination vertex of the source vertex in a graph. The collection of all linknodes “belonging to” some object X is called the chain of X. What all linknodes in a chain have in common is that they share head ID. The next pointer (at the bottom of Figure 2a) is added to allow the GDB to “traverse” (sequentially traverse) a chain, even if its linknodes are not stored sequentially in memory (i.e., the chain is “fragmented”). Starting from the head of a chain, next points to another linknode belonging to the same source vertex, and effectively turns a chain into a traversable linked list. Thus, an object with $N$ properties will feature $N$ linknodes arranged into a linked list. The objective here is to allow a near-memory processor to autonomously “discover” all linknodes attached to a particular source vertex/head ID without having to query the entire memory (e.g., via broadcast), which could be extremely large. Compare the cost of energising 32 billion memory entries to check if they are of interest, to following a couple of hundred linknodes by “hopping” from one to the other using the next pointer. Thus, the implementation of a next pointer is very much rooted in hardware considerations. We have now explained the entire structure of the proto-linknode.
The question now arises: when forming a new chain how should the linknodes be ordered in the linked list and particularly, who should be first in the list? In this version of Views, we define a special linknode flavour called headnode (its primitive “proto” version is shown in Figure 2b) and we do not impose any particular requirements on the ordering of the rest of the linknodes. Intuitively we can think of a headnode as stating that “the object at [link address] exists as an entity”. Practically, it acts as the origin of X’s chain. The headnode has the exact same structure as any other linknode; the difference is in its contents. Within headnodes exclusively: head ID points to its own address (link address) and both primIDs are empty (NULL). The headnode’s self-reference distinguishes it from linknodes as a discrete source vertex. Finally, we note that to end a chain, we fill the final linknode’s next pointer to a specially chosen value that represents the end of chain (EOC). EOC defines the end point of a chain to effectively terminate traversal operations. This is similar to the usage of end-of-file markers in file systems. Therefore, all chains in Views are finite in runtime by the EOC and can be traversed in a linear fashion.
With the ability to create chains, we can now continue building our GDB. We begin by noting that – crucially – in this version of Views primIDs from any linknode point to headnodes. This is how connections are made from a chain to other chains and how a source vertex is connected to a destination vertex via an edge in practice. Other versions of Views may relax that requirement and impose ordering criteria on the formation of the linked list, but these lie beyond the scope of this paper. Examining Figure 2 we note that we now know exactly how to populate all physical fields of any number of headnodes and linknodes that we may have in the database. To illustrate this, Figure 3 shows how a source vertex with 3x (directed) connections maps to a 4-node chain.
0x00ablacknaughtycatcolourtemperspecies
0x00aNULLNULL0x01dtempernaughty0x00bcolourblack0xfedspeciescatEOC
Figure 3: A semantic sentence “Object 0x00a is a naughty black cat” equivalently stored in: (a) a graph where a vertex has degree 3, (b) a Views chain with length 4. Note that the headnode link address oval has been coloured red to highlight that it is a headnode. EOC is a special value used to indicate the end of a chain instead of valid linknodes.
We note that due to the presence of the headnode a vertex of degree $\delta$ maps to a chain of length $\delta$+1:
$l(v)=\delta(v)+1$ (1)
where $\delta(v)$ is the degree of vertex $v$, and $l(v)$ is the length of the linked list of vertex $v$.
2.3 The full Views Linknode: Adding Labellability
link addresshead IDprimID1primID2 prop1 prop2 next
link addresshead ID NULL NULL head properties NULL next
Figure 4: Views GDB model (a) linknode, and (b) headnode taking the same format from Figure 2. prop1 and prop2 are supplemented into our data structure. Note that the properties of the source vertex itself are stored in the location of prop1 in its headnode, to which its head ID points.
We now extend our proto-linknode structure to support recursive properties and become traversable in an elegant way. This upgrades the basic structure described in Figure 1a into what we see in Figure 4a: the full Views linknode data structure. Internally, a linknode is formatted as follows: [“source vertex”, “edge”, “edge properties”, “destination vertex”, “destination vertex properties”, “next linknode”], abbreviated to: [head ID, primID1, prop1, primID2, prop2, next], where primID stands for “primary ID” and “prop” for “properties”, illustrated in Figure 4a. The corresponding structure of headnode format is illustrated in Figure 4b.
In the previous section, we have covered the “labelling” process for source vertices, but edges and destination vertices have not yet been “labelled”. To do so, we furnish the edge and destination vertex entities with relationship-specific properties, illustrated in Figure 4a at the lower left and lower right (prop1 and prop2), to enable recursive labellability. As before, these are also numbers stored in a physical memory and each number corresponds to an address in memory, i.e., they are pointers. They point to linknodes that elaborate on the properties of primID1 and primID2 within the context of the linknode’s triplet, i.e., within the context of the relationship as a whole. The “context-free” versus “context-dependent” distinction is critical: In abstract terms it denotes the difference between “a property of object X” and “a property of object X when it relates to Y via Z”.
In the schema this model was conceived for, a primID is essentially an identifier number as a pointer, but it can point to either an independent object (e.g., “that chestnut brown bird on this green tree”) or a concept (or class, “sparrow” and “camphor tree”). Or in short, a primID refers to a specific entity that is described by its own linknodes. prop pointers then handle context-specific information, which has no meaning of its own outside the specific context of the relation being specified by the linknode. This grants objects the same processing level as classes, so as to avoid the scalability problem of recursive inheritances in GDB expressions. It is thus feasible to build a knowledge graph with uniform data structures, while maintaining powerful contextual descriptions.
CatFelidaefamilytaxonomic ranktype
CatfamilyFelidaeFamilytype taxonomic rank part of speech noun
CatfamilyFelidaetype taxonomic rank Family part of speech noun
Figure 5: Examples of secondary labelling in (a) a traditional directed graph, (b) a Views-based GDB, where the “family” chain has been set up to refer to the taxonomic rank specifically, and (c) another Views-based GDB, where the shown “family” chain has been set up to represent a more generic concept of the term, only specifying the part of speech. Note that the white “family” is nothing but pointer to the head ID of a linked list, within which linknodes define it further.
Note, however, that whilst this is a “natural” choice, it is not obligatory. In particular Views-based GDB instantiations we may want to reserve headnode for classes and treat all individuals as sub-chains. Whether that is sufficient to uniquely identify the individual in question (for example, a particular dog) is a matter for the database contents. This is where a database engineer is needed to determine the specific schema to be used by the database under consideration. The underlying data structure allows the flexibility to choose what will be treated as a headnode.
Illustrative differences can be found between Figure 5b and Figure 5c, both of which describe the nested relationships in Figure 5a: it is up to the database schema whether to treat word meanings as context-dependent properties or as separate linknodes of their own. The linknode on the left (shrunk to a small orange circle for brevity) tells us that “cats belong to the family of Felidae”. In this particular example the “family” primID points to the head of the chain that contains information on what a “family” is, independent of context. Following the link and inspecting the chain we find that this particular “family” chain refers to the taxonomical context and is itself a noun. This does not depend in any way on the fact that cats belong to the family of Felidae, and in fact can be pointed to by multiple primIDs scattered throughout the database.
In contrast, in the alternative example of Figure 5c the primID “family” points to the head of a chain where we have chosen to represent the generic concept of a family. In this example, this “family” chain informs us that it is a noun in a context-independent manner. On the other hand, we note that we have now used prop1 to indicate that in the context of “cats belonging to the family of Felidae”, “family” refers to the taxonomical interpretation. Beyond that point, the chain for “family” may have further specifications as to what the different meanings of the term imply, and the information encoded in the linknode emitted from prop1 can be used to seek that information.
With the above in mind, we note that in recursively labelled graphs, an edge itself can be further linked to other vertices via labelled edges, and so forth [18, 16] (Figure 5a). In Views this is enabled by the “prop” pointers and additionally destination vertices can also be labelled by in-context properties in the exact same manner as edges.
This Soupcontainschickenpartbreastshapecubes marinated in soy sauce Chickenspecies Gallus gallus EOC
Filmis a form of visual storytelling through a sequence of moving images EOC
Figure 6: Examples of: (a) An in-context subordinate chain of length ¿ 1 (“This soup contains chicken, about which we know that [the part is breast, it is in cubes and it is marinated in soy sauce].” and “The species of name for chicken is Gallus gallus”), and (b) Multiple levels of in-context labelling, which reads as Film-is a-form{of-visual storytelling[through-a sequence(of-moving images)]}, or simply “Film is a form of visual storytelling through a sequence of moving images”. Note how we explicitly show pointers to “end of chain” (EOC) to indicate that each green-labelled linknode is not part of a single chain, but rather its own subordinate chain.
In a similar vein to chaining in Section 2.2, successive linknodes can be emitted from prop1 or prop2, forming what we shall call “subordinate chains” (sub-chains). Figure 6a provides an example of a subordinate chain. The example states that a particular soup contains chicken, about which we know that the part of interest is “breast”, and that it is chopped into cubes and marinated in soy sauce.
This information is encoded as a subordinate chain elaborating on the concept of “chicken” in-context. This implements the local infinite labellability for primIDs and aids subtree search and graph mining under the proposed model. [19, 20] We make two further notes: First, a prop pointer points directly to the first linknode of a subordinate chain, and that the linknode that emits the sub-chains acts as the starting node, using the parent context for identification. Second, a sub-chain can emit its own sub-chains, ad infinitum, as shown in Figure 6b. This emphasises the infinitely recursive labellability of the model via recursion into sub-chains of arbitrary depths, and in this way, the proposed GDB model is capable of representing complex contextual information.
2.4 Building a Views-based Graph Database
Tom Hanks0x1act inthis film0xeasSully0x2won2 Oscars0xffor best actor EOCAct In0x4is a cinematic term This Film0x6is afilm0x7title“Sully”0x8protagonistSullyFilm0xais aform Sully Sullenberger 0xcis a public figure 0xdprofessionpilot
Figure 7: A GDB example containing 5 chains: Tom Hanks, Act In, This Film, Sully Sullenberger and Film, interconnected by pointers. Dashed red arrows indicate the traversal path to retrieve the contents from the primID1 pointer at address 0x1. The full contents of the chain Film can be found in Figure 6b.
We are now ready to build a small GDB using the Views model and see its various features in action, as well as introduce some new details, as we shall see. The example GDB is illustrated in Figure 7. Note how in this example each linknode has been explicitly furnished with a physical address, visible inside the orange or green circle that represents it.
First let’s note that the relationship encoding the phrase “Tom Hanks - acts in - this film” is covered by the linknode at address 0x1, where head ID = “Tom Hanks”, primID1 = “this film” and primID2 = “acts in”. Note that “this film” is just some generic, human-readable code-phrase we use to denote the chain that contains information on the film “Sully”; we could have equally well used any other code-phrase such as “linknode at 0x5” or “entity 83”. If we want to retrieve information about the film in question, we follow the pointer from primID1. This leads to the “This Film” chain where we have stored three linknodes (0x6, 0x7 and 0x8). Following the retrieval path indicated by the dashed red arrows, they inform us that: (i) the entity is a film, (ii) the title is “Sully” and (iii) the protagonist is Sully. Similarly, if we want to find out more about the general concept of “act in”, we follow primID2 towards 0x4.
Next, what if we want to find contextual information, such as: “Who in this film does Tom Hanks act as specifically (answer: the character, captain Sully Sullenberger)?” Following primID2 does not lead to an answer; it only returns general and non-contextual information about what it means to “act in” (e.g., it is a “cinematic term”). Thus, we create a subordinate chain and attach the green-coloured linknode at 0xe, specifying that “act(s) in” has the additional property “as - Sully”; the same “Sully” that linknode 0x8 from the chain “This Film” is pointing to. The main question left now is, why not add linknode 0xe directly into the chain of “Tom Hanks”? Because “as - Sully” is a property of “act(s) in” within the context of “This Film”, not of Tom Hanks in general. Note: once again “Sully” is a human-readable codeword. The key feature is that the corresponding primID from the linknode at 0xe points to the “Sully Sullenberger” chain.
We continue exploring the information contained in the database. The node at 0x7 informs us that the chain describes a film titled “Sully”. This is denoted explicitly as “Sully” to show that the pointer does not point to a linknode of the person Sully Sullenberger, but rather to a generic string. This newly introduced feature is key: it allows the GDB to refer to objects outside its direct space of linknodes and thus “ground” itself semantically. In this manner, arbitrary objects such as strings, multimedia, and even ensemble activations of neural networks may be connected to the fabric of the GDB. Further study of this, however, lies outside the scope of this paper.
Next, let us consider the question: “Who is Sully?” If we can identify the chain with headnode “Sully Sullenberger” as the “Sully”; in question, we can read the chain containing linknodes 0xc and 0xd, and find out that he is a public figure and a pilot by profession. Note, however, that in this example reading the chain will not inform us that “Sully” is a protagonist in the film titled “Sully”. This is equivalent to being prompted with “What do you know about Sully Sullenberger?” and replying with the contents of the corresponding chain. The information that he is a protagonist in the film titled “Sully” does not “spring to mind”. To obtain that information one needs a prompt of a sort like: “In what film is Sully Sullenberger a protagonist”? Now we have multiple cues: we know we are looking for an intersection of “Sully Sullenberger” and the concept of “protagonist”, both of whose headnode addresses we assume to know (that is the meaning of being able to “cue” these concepts). Critically, a content-addressable search in the database for where the cued concepts meet will allow us to recover the answer to our question, even though it lies in a linknode that does not exist in the chains of either of the cues! (Instead it is to be found in a third chain, “This Film”.)
Our next case will very briefly exemplify follow-up questions and knowledge build-up. One may ask: “Where does Tom Hanks act?” and receive the answer “He acts in Sully.” Someone unfamiliar with the film may ask: “What is Sully?” and receive the answer “It is a film.” This allows the listener to start building the chain that we have denoted here as “This film” (in fact there is enough information to build the headnode and linknodes at 0x6 and 0x7). By someone not knowing what a film is, the follow-up question “What is a film?” can be also asked, and so on.
For our final example let us consider the following: Suppose we want to make a clear distinction between the real Sully Sullenberger and his on-screen character. In this case we have two immediately obvious options: either (i) we will start a subordinate chain for the primID pointing to “Sully” in the linknode at 0x8, or (ii) we will create a new chain for the dramatised version of Sully Sullenberger and let (or indeed “rewire”) the connection from the primID of the linknode at 0x8 to this new chain, and then somewhere in the new chain introduce a linknode making it clear that the chain refers to the dramatis personae of the real Sully Sullenberger. This illustrates both the importance of setting up a schema so as to be fit-for-purpose and the flexibility of the Views data structure. We note that the specific mechanics allowing “rewiring” as illustrated here bear a more than passing resemblance to the notion of schema learning in psychology, but further exploration of this concept lies outside the scope of this paper.
We conclude the section with the remark that Views is also naturally compatible with the property graph data model with “nodes” mapping to “headnodes”, “edges” mapping to primIDs and “properties” possible to attach via subordinate chains.
3 Hardware Implementation
The proposed model was designed with hardware-friendliness in terms of storage and performance in mind. This grants the model a high degree of storage efficiency and response speed. (See examples in Section 4.) Furthermore, the structure of linknodes strongly prescribes implementations where each element of the linknode is stored in a separate memory array.
In this section, we will first present two possible mappings of the Views data structure into hardware; two ways to map linknode elements to physical memory arrays: the “CNSM” and the “normalised” mappings. Then, we will briefly introduce the ASOCA implementation for hardware acceleration of GDBs in order to: a) illustrate the hardware-friendliness of the approach and relative simplicity of implementation and b) allow the reader to follow the discussion in section 4 with a better understanding of how the hardware works “under the hood”.
3.1 Mapping linknodes to physical arrays: Two Allocations
Eight functionally identical memory arrays are used in the CNSM allocation. Each memory array is allocated an identifier reflecting its functional meaning as shown in Table 1. “C” arrays store primIDs, i.e., the main “Content” of the Views-based GDB. “N” arrays store pointers that allow traversals, i.e., they are “Navigational” in nature. “S” arrays store pointers that branch towards “Subordinates”. Finally, “M” arrays were added to store extra properties (“Miscellaneous”) while bringing the total number of arrays to a power of 2. This implementation is used in the demonstration of Section 4.
Table 1: CNSM Allocation by Array
Type/function Identifier Linknode mapping Usage
Content C1 primID1 Edge vertex pointer
C2 primID2 Destination vertex pointer
Navigator N1 head ID Source vertex pointer
N2 next Next linknode pointer
Subordinate S1 prop1 Edge subordinate
S2 prop2 Destination subordinate
Miscellaneous M1 prop1 Edge universal properties
M2 prop2 Destination universal properties
We now turn to M arrays. They are introduced for extra convenience when translating the Views data structure into hardware, allowing storage of properties that every graph edge or vertex can be expected to have in common schemas, such as edge weights, vertex degrees and quantifiers for 1st order logic among many other possibilities. For brevity, we will refer to these as “universals”. The configuration of universals is an implementation-specific optimisation for fast in-situ property information retrieval, very much dependent on the database engineer’s needs. Notably, universals could be stored as additional linknodes, but their universal applicability means they can be essentially “hard-wired” within the M arrays.
Finally, we note that even schemas such as CNSM can be further tweaked. In this particular example we chose to designate C1 as “edge pointer” and C2 as “destination vertex pointer”, clearly segregating edges from destination vertices. This is not obligatory, however, as pointed out in the previous section.
Table 2: Normalised Allocation by Array
Type Identifier Linknode mapping Usage
Content C1 primID1 Edge vertex pointer
C2 primID2 Destination vertex pointer
Navigator N1 head ID Source vertex pointer
N2 next Next linknode pointer
Next, the Normalised allocation is a minimalistic version for simpler databases, emphasising compactness of representation over complete functional flexibility. As illustrated in Table 2, S and M arrays are removed, leaving just C and N behind. This allocation may still represent subordinate chains by treating them as separate chains, but analysis of this possibility lies outside the scope of this paper. The normalised allocation is more suitable for graphs with less context-dependent information. Finally, we note that S arrays and/or M arrays can be optionally supplemented upon the Normalised allocation to produce further options for the GDB designer.
3.2 The ASOCA implementation
Our Associative Chip Architecture (ASOCA) aims to turn the ideas behind Views into a series of hardware accelerators for GDB operations. We started by implementing a memory array; the Associative Memory Chip I (ASOCA1). Using beyond-CMOS [21] memory cells as the basic bit-storage unit, it is a dually-addressable memory (DAM) array capable of storing any of the pointers in the CNSM and Normalised schemes shown previously (i.e., any line in Table 1 and Table 2) and promises good power performance (further details in [22, 23]). ASOCA1 arrays are designed to hold $64\times 64$-bit pointer entries each and for the purposes of this paper, they can be considered the “unit storage array”.
[…]
With the basics of the hardware implementation presented, we now cover the corresponding instruction set architecture (ISA), i.e., the set of assembly-level operations required for operating on graph-structured data. In this paper, we will restrict ourselves to only presenting key operations at the conceptual level. We will use these in the discussion of the next section.
The ASOCA2 non-trivial operations are as follows:
  1. 1.
    Program (PROG): Sets a pointer within a linknode (e.g. the primID1 pointer of some linknode N, stored in a C1-type array within a supercluster).
  2. 2.
    Address-addressable read (AAR): Reads the pointer stored at a specified address in the database. It acts as a standard memory read in conventional memories.
  3. 3.
    Content-addressable read (CAR): A pointer is searched for in the database as a cue. The address(es) where it has been found are returned by the underlying associative memory. We can specify which array within a supercluster to search (e.g. in CNSM mapping: C1, C2, N1, …).
  4. 4.
[…]
  5. 5.
    A set of extra utility operations (HEAD, CARNEXT & TAIL): Accelerated composite operations that allow the hardware to efficiently traverse the graph structure. HEAD reads N1 of a given linknode and finds the headnode of the chain that “owns” this linknode. CARNEXT returns “the next match” in the event that a CAR/CAR2 operation identifies more than a single match/answer. TAIL iteratively reads N2 until the EOC and returns the address of the last linknode of the chain that ‘owns’ a given linknode.
Here is a basic example of how the ISA can be used to answer simple queries: after programming a complete database into the hardware with PROGs, massively parallel CARs on N1-type arrays can be used to quickly find the addresses of all linknodes belonging to a specific headnode (i.e., they can “highlight” a complete chain). This can be used directly for queries of the sort: “Fetch all information directly associated with Tom Hanks.” Note the use of “directly” to denote that we are not looking for information about Tom Hanks available in other chains, i.e. cases where thinking of a film or some other concept would “make us think of” Tom Hanks.
Afterwards, AARs can be used to retrieve further contents from the identified linknodes (e.g. the primIDs) and fuel further searches using CAR, AAR, HEAD, CARNEXT or TAIL. Another example would be using CAR2 to locate 2 out of the 3 components in a ternary relationship, using a follow-up AAR to retrieve the final component. This can answer queries like: “Who won 2 Oscars?” (see Figure 7). We would send a CAR2 operation using “won” and “2 Oscars” as the query and then trace the head of the chain that owns nodes that match that description.
The ISA under discussion is natively powered by Views to support the scalability and heterogeneity of data storage, while associative search ability from hardware memory arrays backs efficient data retrieval from Views-based GDBs. It denotes the strong synergy between the Views data structure and the ASOCA hardware architecture. Beyond advanced graph operations such as subgraph matching and graph traversal, this co-design methodology provides the feasibility for the dedicated hardware acceleration of optimised graph algorithms by utilising both the GDB model and the hardware architecture [25, 26].
3.3 Hardware Storage Comparisons
We compare the storage performance of a Views-based system against 4 representative GDB systems: Neo4j (LPG), Memgraph (LPG), Apache Jena TDB2(RDF) and Blazegraph (RDF). The benchmark reproduces the minimal dataset of Figure 7 in each implementation and in Views, and then measures total on-disk footprint.
Within LPG implementations, Neo4j uses native on-disk graph storage across multiple files, often fixed-size. Its disk usage is calculated from entity/property counts and documented per-record overheads [27]. Memgraph is primarily configured as an in-memory graph database, and we report its write-ahead log for disk usage.
Conventional RDF stores serialise datasets into plain formats for on-disk storage [14, 11], while both Blazegraph and Apache Jena TDB2 back indexes in B+ tree files. Blazegraph keeps a single journal file, and we extract its disk usage from internal statistics. TDB2 employs multiple files for indexing and doesn’t scale well with smaller datasets; we report the decompressed size of its N-Quads snapshot instead for comparison.
For Views, we implement the CNSM allocation on one ASOCA2 supercluster and measure the occupied memory after programming the dataset. Oversized strings are outsourced and indexed by primIDs; their sizes are included.
For reference purpose, we also record here each database’s dump size where possible (e.g., a snapshot or backup). For Views, this is the total disk usage with entry addresses included. The results are summarised in Table 3 and Figure 9.
Table 3: Storage Performance Comparison across Views, RDF and LPG Implementations
Database Disk usage (Bytes) Entity number Dump size (Bytes)
Neo4j 1554 24 7521
Memgraph 7362 24 446 (compressed)
Blazegraph 11697 39
Apache Jena TDB2 4014 39 540 (compressed)
Views 685 19 756
Figure 9: Storage efficiency comparison among different graph database implementations: (a) Disk usage in Bytes; (b) Number of entities stored.
Views stores relationships as directly addressable linknodes with no record headers or secondary indexes, which removes per-entity overhead as seen in RDF or LPGs. Second, pointer-based sharing avoids duplication: Views collapses a triple into one linknode plus reusable headnodes, while labels and objects are all referenced via headnodes. Last but not least, context-specific information is stored in sub-chains to enable direct hardware support for data locality and performance, rather than separate triples or property rows.
As a result, Views’ linknode data structure together with the matching hardware implementation means the model keeps bytes-per-relationship essentially flat, preserving its significant advantage in storage efficiency, retrieval performance and scalability. Still, note that some database implementations are affected by their on-disk storage format and page alignment. Optimisations, e.g., compression or index layout choices, can improve the numbers of existing GDB implementations quoted in Table 3. Views does not seek to entirely replace them, but instead offers strong compatibility with them as the underlying GDB model.
4 Operation Examples
In this section, we demonstrate some more elaborate, more practical applications of the proposed GDB model, highlighting its capabilities in semantic reasoning and cognitive modelling through examples and analysis.
4.1 Semantic Reasoning
Vector symbolic architectures (VSA), also known as hyperdimensional computing, show great potential in solving the compositional problem in traditional deep learning [28, 29, 30, 31, 32]. They achieve this by supporting “semantic reasoning”, i.e. representing semantically meaningful concepts as vectors and then performing “clean”, symbol-level manipulation upon them [33, 34]. Frequently, semantic reasoning is underpinned by a vector database; however, here we show how the Views model can use GDB-oriented methods to support this functionality. To illustrate this we examine a typical syllogistic example in natural language whereby we wish to deduce that “‘This’ (i.e. the object of discourse) is feline”:
Major Premise: ‘This’ is a cat;
Minor Premise: Cats are feline;
Conclusion: ‘This’ is feline.
[…]
0x6fdCatfamilyFelidae NULL NULL EOC
Figure 10: Part of the example knowledge base contents: (a) Chains of 0x00a (from Figure 3b) Black and Cat in Views format. (b) The family-Felidae linknode (shown above in red, dashed lines in the Cat chain), denoting the link among cat, Felidae and family.
To achieve this, we need to represent the knowledge encoded in the premises and write a small programme that employs calls to our Views-based GDB to work through the necessary logical steps for deriving the conclusion. For the representation let us choose a straightforward scheme: First, a chain to represent the object we are attempting to make an inference about (“This”) including a linknode with primIDs “species” and “cat” (just as we have done in Figure 3b). Then, a chain for the general concept of a “cat” featuring a linknode whose primIDs are the pair (“family”, “Felidae”) plus any other relevant chains. These are illustrated in Figure 10a.
Next, comes the programme. This can be code running on any host machine with access to a Views-based GDB. We will focus on the calls to the GDB here before presenting the full pseudocode in Algorithm 1. Throughout this example, we write the database call code in a form that closely resembles Python for simplicity and familiarity. We also omit trivial checks such as verifying that our CAR/AAR operations return valid results to focus on the logic of the algorithm rather than the minutiae.
In our example solution, we begin by effectively asking our database: “What family does ‘this’ belong to?” In terms of database instructions, the system issues a pair of CAR2 operations:
    results_C1 = CAR2(N1="this", C1="family")
    results_C2 = CAR2(N1="this", C2="family")
    results = [results_C1, results_C2]
where our notation assumes CNSM configuration. The main code can then check if the results list contains the (“family”, “Felidae”) pairing. If yes, the query has been answered. However, from the way we constructed the example, we know that this is not the case and therefore another stage of reasoning is required.
In the second stage, we will attempt to find the species of ‘this’ and then see if we can infer the family from the species. We therefore perform the following queries to the database:
    addresses_1 = CAR2(N1="this", C1="species")
    addresses_2 = CAR2(N1="this", C2="species")
    results_C2 = AAR(addresses_1, C2)
    results_C1 = AAR(addresses_2, C1)
    results = [results_C1, results_C2]
The CAR2 pair informs us whether ‘this’ is in any way related to “species” and stores the addresses of the corresponding linknodes, from which we again perform AARs to tell us how ‘this’ relates to “species”. For each result in results (here the headnode of “Cat”) we now ask the question: “Is the species under discussion a member of the family of Felidae?”, or in database query terms:
    results_C1 = CAR2(N1=results, C1="family")
    results_C2 = CAR2(N1=results, C2="family")
    results = [results_C1, results_C2]
With every set of updated results the programme can check for the pairing (“family”, “Felidae”) and once it is found, the query has been answered in the affirmative. Notably, other programs that solve the task (in this representation scheme) are also possible.
The full pseudocode for this search process is illustrated in Algorithm 1 where operations calling on our Views-based GDB have been highlighted in red. Note how we compact the pairs on C1/C2 into a single pseudocode function for brevity.
Algorithm 1 Search for “Felidae” in ‘this’ chain and its species chain.
[…]
2:felidaeAddrfelidaeAddr
3:for thisAddrthisAddr in ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}CAR2}(N1=\texttt{0x00a},C1/C2=\texttt{family})$ do
4:  $thisResult\leftarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}AAR}(thisAddr,C1/C2)$
5:  if ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}AAR}(thisAddr,C2/C1)=\texttt{Felidae}$ then
6:   return thisAddrthisAddr $\triangleright$ Felidae found in ‘this’ chain
7:  end if
8:end for
9:for thisAddrthisAddr in ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}CAR2}(N1=\texttt{0x00a},C1/C2=\texttt{species})$ do
10:  $thisResult\leftarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}AAR}(thisAddr,C2/C1)$ $\triangleright$ Species of ‘this’
11:  for $speciesAddr$ in ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}CAR2}(N1=thisResult,C1/C2=\texttt{species})$ do
12:   $speciesResult\leftarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}AAR}(speciesAddr,C1/C2)$
13:   if ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}AAR}(speciesAddr,C2/C1)=\texttt{Felidae}$ then
14:     return speciesAddrspeciesAddr $\triangleright$ Felidae found in its species chain
15:   end if
16:  end for
17:end for
18:return NULLNULL $\triangleright$ Felidae NOT found
We finish the discussion of this example by noting that different data representation schemes can change the logic of the solution very dramatically. Consider, for example, how the programme would change if instead of encoding the statement that “The species of ‘this’ is cat” as illustrated in Figure 10 we encoded it with the primID pairing (“is”, “cat”). Or imagine in other very hierarchically structured taxonomical database(s), information about “family” can only be traced by its subordinate “genus”. This illustrates the depth and diversity of how Views can be used. Interested readers are invited to develop creatively their own bespoke examples.
4.2 Cognitive Processing Application
Copycat is a cognitive model imitating human analogy, which is designed to answer string analogy problems such as “if $abc\sim abz$, then $zyx\sim?$” or “abc:abz::zyx:?abc:abz::zyx:?[17, 35]. The model uses a static concept storage structure, organised as a graph consisting of vertices representing “crisp” (i.e. not probabilistic or otherwise “fuzzy”) concepts. This was called the slipnet.
The name “slipnet” comes from the key operating principle of Copycat, the slippage. It is a mechanism whereby a concept can be dynamically substituted by another during problem-solving [36]. For example, in the string analogy example above, Copycat may test the hypothesis that [3rd-letter-in-string is last-letter-of-alphabet]. However, an alternative hypothesis is that [3rd-letter-in-string is first-letter-of-alphabet], and in order for Copycat to “think” of this possibility, the model employs slippage: It replaces “Last” with “First”; a pair of concepts that are connected in the slipnet by an edge labelled ”Opposite”, as illustrated in Figure 11. As such, slippage allows Copycat to flexibly explore alternative hypotheses, even when concepts appear semantically inconsistent or contradictory, e.g., First-Last or Leftmost-Rightmost.
0xcafeFirstoppositelast slip lock=True slip lock=False 0xcaffLast0xc0c0 Activ=30<80, Activ lock=True Opposite0xface Activ=100>80, Activ lock=False Activ propagatingSlip! (Slipnet threshold is 80)
Figure 11: Example of slippage from Last to First via Opposite in Views slipnet. The example database contains a very simple proposition, saying “First is the opposite of last” at address 0xcafe. Note the purple rectangles representing the contents of the M arrays. The slippage process begins with the activation propagating from First to Opposite, allowed because Activ lock is false (blue path). As a result, the activation value of Opposite crosses the threshold (80), which then triggers the slippage to First via Opposite, after a slip-lock check conducted at Last (0xc0c0).
Here, we detail the representation of the slipnet using Views, along with the implementation of the slippage mechanism. Table 4 summarises slipnet’s data organisation converted into ViewsCNSM allocation. This conversion of slipnet starts by storing each concept (known as slipnodes in Copycat’s terminology) as a headnode in the N1 array (e.g. “First”). Once slipnodes are instantiated, conceptual relationships are encoded using the regular (non-head) linknode format: C1 holds the primID pointer to the edge label, and C2 holds another to the destination concept. This directly implements the sliplinks, such as “First – Opposite – Last”, within the Views framework. Note that some edges in Copycat are not explicitly labelled, in which case we assign said edges ad hoc IDs pointing to headnodes without further linknodes in their chains.
Copycat is also an “activation-based” cognitive model and as with other such models [37, 38, 39, 40], it works by propagating and periodically updating the levels of activation across the slipnet following predefined operation rules [17]. This mechanism affects how operators in this model seek both regular and creative (slipped) connections, mirroring the cognitive processes underlying human creativity. In order to operate Copycat-like activation-based mechanics under Views we need to allocate space for storing: a) the activation value/strength “Activ” of each linknode, b) a “natural activation decay” value regulating how quickly activation decays in the absence of other updates (in original Copycat: “conceptual depth”; for details please see original publication [17]) and c) a “conductance” value encoding how efficiently activation can spread over a linknode. We also add, of our own accord, a couple of 1-bit flags for improving the operational flexibility of the system: the slip lock which enables/disables slippage along a given primID and the activation lock (“Activ lock”) which enables/disables changes to “Activ”. We define these as scalars or boolean as appropriate and allocate them to the M1/M2 as universal properties, shown in Table 4. Finally, for simplicity we will use primID1 to denote edges and primID2 to denote destination vertices.
Table 4: Slipnet data under CNSM allocation. Items in bold denote data objects used for running Copycat-like activation dynamics. Note, for example, how destination vertex is represented as a pointer and stored in array C2 (making it a primID). Similarly, note how the concept of slip lock is represented as a bool and stored within either M1 or M2 arrays.
Content Type Linknode mapping Array
source vertex pointer head ID N1
Next linknode address next N2
edge label pointer primID1 C1
destination vertex pointer primID2 C2
Subordinate I reserved prop1 S1
Subordinate II reserved prop2 S2
Conceptual depth scalar prop1 M1 (headnode)
Activ scalar prop1 M1 (headnode)
Activ lock bool prop1 M1 (headnode)
Conductance scalar prop1/prop2 M1/M2 (1/linknode)
Slip lock bool prop1/prop2 M1/M2 (1/primID)
With all the data needed to run activation mechanics now represented, we can set up a simple example of how activation might work. For an iteration of the activation propagation from the source vertex to the destination vertex, we execute the following pseudocode in the current linknode:
    if primID1.actiLock == 0:
        primID1.activ = primID1.activ * primID1.conceptualDepth
                        + currLink.head.activ * currLink.conductance
The code effectively states that “If the edge headnode is not activation-locked, its activation level will be updated as follows: it will decay by a factor ($\leq 1$) determined by its conceptual depth and increase by a fraction ($\geq 0$) of the activation of the headnode hosting the chain of the current linknode, where the fraction is determined by the conductance of the active linknode”. Then we continue with:
    if (primID1.activ > slipnet.threshold) and (primID2.slipLock == 0):
        currLink.head.slippingFrom.append(primID2.head)
which checks whether the activation of the edge headnode (primID1) is greater than a preset threshold value, and whether the edge to the destination vertex (via primID2) is not slip-locked. If both conditions are met, the destination vertex is added to the list of slippage candidates (currLink.head.slippingFrom) of the source vertex. We note how assigning slip locks to each individual linknode can selectively activate/deactivate slippage along individual edges in the slipnet’s graph. We further note that far more elaborate programmes can be written for implementing slippage.
We close this section by noting that the example above illustrates how the Views format supports non-trivial cognitive applications and offers tantalising hints towards how the hardware described in Section 3.2 may be extended to accelerate the operations connected to activation and slippage management. Finally, as a matter of interest, the slipnet of the original Copycat from [35] transposed into Views format results in 77 headnodes across 11 categories, interconnected by 195 linknodes.
5 Discussion
By explicitly supporting the storage of semantic triples, the Views GDB model joins a broad community of graph representations, such as RDF triples and edge lists, while introducing distinctive capabilities (so it can be considered as a “triples+” kind of representation). It would be an interesting direction of future investigation to build a comprehensive table of mappings between different existing representations and Views, i.e. to determine the Views schemas to be used in order to emulate said representations. For example, there already exists literature that shows how edge list-based representations can be transformed into adjacency list-based ones [41], where Views can be regarded as an example of the latter. This bodes well for portability via automated translation between formats.
Tom Hanks 0x1 this film nil act in 0xe nil as Sully nil 0x2 nil won nil 2 Oscars 0xf nil for best actor nil
Tom Hanks0x1act inthis film0xeasSully0x2won2 Oscars0xffor best actor EOC
Figure 12: Example of (a) a Lisp cons structure representing (b) the corresponding Views Tom Hanks chain. Addresses are set and coloured accordingly to demonstrate their equivalence. Nested linknodes form subordinate chains, illustrating hierarchical relationships (e.g., “2 Oscars for best actor”). Note that in the cons structure, the car fields above act in and 2 Oscars correspond to entry points to local primIDs and their respective subordinate chains inViews. nil terminators mark the ends of chains, preserving list structure integrity. Naturally, other configurations to represent the information in both formats are possible.
Another example of portability concerns list-based structures. The fundamental abstractions of linked lists, as in Lisp’s cons, car, and cdr primitives [42], have long served as a foundation for representing hierarchical and relational data in symbolic computation. This aligns with the organisation of Views, where node relationships and traversal logic take the form of linked-list navigation. In Views a linknode can act as a Lisp cons cell: one of the primIDs (for example primID1) acts as car, whilst the next pointer, pointing to the next cons cell within the linked list, acts as the cdr. Additionally, each linknode “equips” the cons structure with 3x additional pointers: primID2 and prop1 and prop2. Finally, NULL is analogous to nil in the Lisp world. In this framework, a Views linknode corresponds to a Lisp cons cell: it pairs a sublist with a next pointer to the subsequent linknode (another cons cell) in a list. Thus, a Views chain acts as a sequence of “Views-enhanced” cons elements, with the added functionality that in Views the primID2, prop1 and prop2 pointers can spawn additional lists.
For example, the “Tom Hanks” chain in Figure 7 can be expressed in Lisp cons and linknode structures in Figure 12 for comparison. The example illustrates how the Views format can be understood under different prisms when considered in relation to well-established representations.
We close this section by noting that beyond static data representations there are tantalising indications that representations of processes might be possible to integrate into the Views GDB model, drawing inspiration from $\lambda$ calculus. This would extend Views’ ability to represent and manipulate procedural constructs, a significant step towards bridging the gap between procedural and declarative knowledge in cognitive architectures [43, 44, 45, 46]. This could prove to be a very interesting line of further investigation.
6 Conclusions
In this paper, we have proposed a GDB model named Views, which supports the notion of graphs with infinitely recursive labellability (whereby edges and vertices, and their properties, and their properties ad infinitum can be labelled) in an intuitive manner, and converts the infinitely recursively labellable graphs into linked list structures for hardware-friendly data storage and graph traversal. Furthermore, we have laid the foundations of a natural hardware implementation of our model and provided a fundamental instruction set architecture (ISA) that can operate on the Views data structure, including massively-parallel content-addressable read operations. Next, we evaluated this hardware implementation’s storage efficiency across existing RDF and LPG-based GDB implementations revealing an advantage in storage efficiency and scalability stemming from the combination of a uniform data structure, a linked-list organisation and tight co-design with the corresponding hardware. This hints towards the potential for tremendous acceleration of graph analytics workloads via parallelisation. Next, we showed examples of how the structure can be used to carry out example reasoning tasks and how it may be naturally extended to accommodate the operational requirements of a non-trivial example cognitive model, Copycat, which uses bespoke features such as its “slippage” mechanism. This shows promise towards eventually underpinning key cognitive tasks such as semantic reasoning, logical deduction, and cognitive processes such as analogy. Finally, we have illustrated how the approach is versatile and highly compatible with (i.e. “admits mappings from”) a range of conventional knowledge graph representations, which indicates a relatively low barrier to the translation from well-established representations into Views. We hope that our proposed data structure and the corresponding hardware implementation outline act as an extra bridge between the computer science & AI community and the hardware design community, inspiring further innovation in hardware-aware AI long into the future.
Acknowledgements
This study was supported by the Engineering and Physical Sciences Research Council (EPSRC) under Grant EP/V008242/2 and Scottish Enterprise High Growth Spinout Programme (HGSP) Atlas, funding reference number PS7305129O.